The height of an acorn falling out of a tree is given by h=-16t squared+b. If an acorn takes 1 second to fall to the ground. What is the value of b?
Answer by nerdybill(7384)

The height of an acorn falling out of a tree is given by h=-16t squared+b. If an acorn takes 1 second to fall to the ground. What is the value of b?
.
If it hits the ground
h = 0
and they give you that
t = 1
.
plug it all into:
h = -16t^2+b
0 = -16(1)^2+b
0 = -16(1)+b
0 = -16+b
16 = b